A 225m long, 300mmϕ concrete pipe and a 400m long, 400mmϕ concrete pipe are connected in series. What is the diameter of a 625m long equivalent pipe?


1
Expert's answer
2022-03-10T10:18:10-0500

Explanations & Calculations


  • If the equivalent diameter is "\small D," then

"\qquad\qquad
\begin{aligned}
\small \frac{L}{D^5}&=\small \Sigma_{i=1}^2\frac{L_i}{D_i^5}\\
\small \frac{625}{D^5}&=\small \frac{225}{300^5}+\frac{400}{400^5}\\
\small D^5&=\small \frac{625}{1.317}\times100^5\\
\small D&=\small 342.97\\
&\approx\small 343\,mm
\end{aligned}"